# Human limits in Machine Learning: Prediction of plant phenotypes using soil microbiome data

### v1

@@ D.2.2 Using Network Comparison for Feature Selection
We use the SPRING method (semi-parametric rank-based approach for inference in graphical model) [33] in the NetComi package to construct two microbial networks: one corresponding to label 00 and one corresponding to label $1$. To build these networks, the estimated partial correlations are transformed into dissimilarities via the signed distance metric and the corresponding similarities are then used as edge weights. We then compare these networks in order to identify OTUs that have the greatest difference based on degree values. Figure 10 shows an example of two microbial networks with different graphical structures, and Table S3 in the Appendix lists OTUs ranked by the difference in degree in the two microbial networks for diseased and non-diseased classes of pitted scab response. The OTUs are sorted based on the values of degree difference column and similar to ML strategy top 30% of them are selected as important features.
D.2.3 Combination of ML and Network Strategies in Feature Selection
We define a scoring value for each OTU based on whether they are identified as important by ML strategy ($score=1$), by the network comparison ($score=2$), or both ($score=3$). If the OTU is not identified as important by any strategy, it is denoted $score=0$. Figure 11 displays the scoring values of all OTUs at the Phylum level for all responses.
D.3 Model descriptions: random forest and Bayesian neural network
We apply two types of classification models to predict potato disease and yield: random forest (RF) and Bayesian neural networks (BNN). We consider separately three types of predictors: OTU abundances (5 taxonomic levels and 20 normalization/zero replacement strategies described in Section D.1.2), environmental predictors such as soil characteristics and microbial population density, and a combination of both types. Table 3 lists all the models we consider in this study.
@@ D.3 Model descriptions: random forest and Bayesian neural network
Table 4 shows the number of predictors included in each model as well as the data choices related to taxonomic level or normalization and zero replacement strategies. For example, using all OTUs (first row), we have five taxonomic levels and 20 normalization+zero replacement strategies each, so in total, we have 100 (20 normalization strategies times 5 taxonomic levels) different OTU datasets. For each of these 100 datasets, the number of predictors (i.e., OTUs) would depend on the taxonomic level being analyzed. For example, at the Phylum level, there might be 42 predictors, while at the Genus level, there could be up to 485 predictors.
ALL-OTU 5 20 42–485
OTU-S0 5 1 7–54
OTU-S1 5 1 5–134
OTU-S2 5 1 5–134
OTU-S3 5 1 7–54
Alpha 5 6 9
Soil 6 12
DS 6 4
D.3.1 Random Forest Model
@@
Here, we set different values for parameters (see Table S4 in the Appendix) and tune them using GridsearchCV to find the optimal values for the RF classifier. GridSearchCV uses a “score” method for evaluating the performance of the cross-validated model on the test set. We use the F1 score to evaluate the performance of the model. The F1 score is the harmonic mean of precision and recall, where the proportional contributions of precision and recall to the F1 score are equal. A F1 score of 1 is the best, while a score of 0 is the worst. Finally, when the parameters for the RF model are tuned, we use 20% of samples to report the performance of the final model. We use the weighted average F1 score to evaluate the performance of the models, which is computed by averaging all the per-class F1 scores while accounting for the number of samples in each class.
D.3.2 Bayesian Neural Network Model
Given that neural network models are intrinsically data-hungry and microbiome applications have relatively small sample sizes, BNN models provide a suitable alternative for small sample sizes as they provide natural protection against overfitting. This is due to the fact that distributions are considered for the parameters in the model which allow us to marginalize them so that the prediction is based on data points alone [9].
We provide details on the BNN model formulation next. Let $\mathcal{D}=\{x_{i},y_{i}\}_{i=1}^{n}$ be the data with $n$ samples where $x_{i}$ is a $N$-dimensional feature vector in $\mathbb{R}^{N}$ and $y_{i}\in\mathbb{R}$ is the target response. A BNN model has $L$ layers with $K_{l}$ neurons in the $l^{th}$ layer. The weight parameters are defined in a set of matrices $\mathcal{W}=\{W_{l}\}_{l=1}^{L}$ where each $W_{l}$ corresponds to the weights for the $l^{th}$ layer, and is of size $(K_{l-1}+1)\times K_{l}$. For an input $x_{i}\in\mathbb{R}^{N}$, the neural network maps it to a response $f(\mathcal{W},\textbf{x})$ by multiplying the input of each layer by the weights and then transforming via an activation function such as rectified linear unit (ReLU): $h(a)=\max(0,a)$.
Unlike regular neural network models, a BNN imposes a prior distribution on the weights $\mathcal{W}\sim p(\mathcal{W})$ to capture their uncertainty. The posterior distribution of the weights given the data is then given by:
$p(\mathcal{W}|\mathcal{D})=\frac{p(\mathcal{W})\prod_{i=1}^{n}p(y_{i}|f(\mathcal{W},x_{i}))}{p(\mathcal{D})}$
@@ D.3.2 Bayesian Neural Network Model
We use the original implementation of BNNs denoted “software for flexible Bayesian modeling and Markov chain sampling" [20] for training and testing on our data. For each BNN model, we use five hidden layers, one input layer and one output layer. The input layer has $n$ input neurons with $n$ being the number of predictors for the model. This number varies depending on the predictors to include. For example, for the ALL-OTU model for the Phylum level, we have 42 predictors (Table 2), and thus the input layer has $N=42$ neurons. For each hidden layer, we have $3N$ hidden neurons and one bias node. For the output layer, we only have one neuron with a binary value for the response. We use the hyperbolic tangent as the activation function: $f(x)=tanh(x)$.
We choose a hierarchical zero-mean Gaussian prior on all weights and on bias nodes. The prior for the weight from node $i$ of one layer to node $j$ of another layer is formulated as follows [20]:
$\mathcal{P}(w_{ij}|\sigma_{w_{i}},\sigma_{a_{j}})=\frac{1}{\sqrt{2\pi}\sigma_{w_{i}}\sigma_{a_{j}}}\exp(-\frac{w_{ij}^{2}}{2\sigma_{w_{i}}^{2}\sigma_{a_{j}}^{2}})$
where $\sigma_{w_{i}}$ are the hyperparameters for the standard deviation of the priors for node $i$ of layer $w$, and $\sigma_{a_{j}}$ is the hyper parameter for the standard deviation of the priors for node $j$ of layer $a$. Let $\tau_{w_{i}}=\sigma_{w_{i}}^{-2}$, then the hyper prior is given by
@@
Given that HMC does not scale well for high dimensional parameter spaces and large datasets [31], we did not fit the BNN on all the models described in Table 3. In particular, we do not fit a BNN model on Genus level for the ALL-OTUs as this model would involve over 9 million weights in the network with 485 input neurons.
D.4 Predictive Power of Microbiome Data: Comparison to Random Data
Strategy 1:
Strategy 2:
Strategy 3:
Real microbiome matrix is used with a permuted response vector.
Strategy 4:
For each strategy, we generate $N=200$ random datasets to train the RF model. Let $F_{i}$ for $i=1,...,N$, denote the weighted F1 scores for the randomized data and let $F_{original}$ be the weighted F1 score on the real data. Let $X=\sum_{i=1}^{N}\mathbb{I}(F_{i}>F_{original})$ be the number of random datasets that perform better than the read data. We use the exceeding value (EV) $EV=\frac{X}{N}$ as test statistic to test the null hypothesis that the real data performs just as random data in the prediction on disease outcomes. The EV denotes the percentage of random datasets that perform better than the real microbiome data out of $N$ random datasets. If $EV<\alpha$, we reject the null hypothesis and conclude that there is predictive power (beyond random) on the real microbiome data. Here, we set $\alpha=0.05$ as in [1].
D.5 Full Model Selection
Full model selection (FMS) [25, 26, 24] involves the process of listing all data preprocessing steps, model options and selection of predictors, and using a decision tree model to identify the choices that yield the highest measure of performance. Here, we fit a FMS strategy with the following options: 1) type of normalization, 2) type of zero replacement, 3) taxonomic level, and 4) data augmentation. We combine the type of normalization and type of zero replacement strategy into one variable (denoted $\text{NM}_{i}$ for $i=1,\dots,20$). We focus on the weighted F1 score as measure of performance, and we include all OTU predictors (that is, we do not consider feature selection as one of the options to compare). We build the regression decision tree by using the DecisionTreeRegressor which is available in scikit-learn [21]. We use the default parameters in the DecisionTreeRegressor such as "squared error" as the criterion to measure the quality of a split, a minimum number of 2 for the samples required to split an internal node, and a minimum number of 1 sample required to be at a leaf node. In order to create an informative decision tree that can be interpreted, we use a maximum depth of 4.
@@ Acknowledgements
[…]
@@
[…]
@@
Figure 3: Boxplots of weighted F1 scores for random forest models for different types of predictors (rows) and different yield or disease outcomes (columns). For description of the rows, see Table 3 in the main text. The range of each boxplot depicts the weighted F1 scores for datasets in different taxonomic levels and different normalization and zero replacement strategies. The dashed line corresponds to the weighted F1 score when fitting the model with random datasets(see Section B.2.1).
Figure 4: Boxplots of weighted F1 scores for Bayesian neural network models for different types of predictors (rows) and different yield or disease outcomes (columns). For description of the rows, see Table 3 in the main text. The range of each boxplot depicts the weighted F1 scores for datasets in different taxonomic levels and different normalization and zero replacement strategies. The dashed line corresponds to the weighted F1 score when fitting the model with random datasets (see Section B.2.1).
@@
Figure 26: Boxplots of the weighted F1 scores by Random Forest (RF) and Bayesian Neural Network (BNN) models for yield by plant using environmental predictors.
Figure 27: Boxplots of the weighted F1 scores by Random Forest (RF) and Bayesian Neural Network (BNN) models for yield by meter using environmental predictors.
Figure 28: Boxplots of the weighted F1 scores by Random Forest (RF) and Bayesian Neural Network (BNN) models for black scurf disease using environmental predictors.
@@
Figure 29: Boxplots of the weighted F1 scores by Random Forest (RF) and Bayesian Neural Network (BNN) models for scab disease outcome using environmental predictors.
@@
Figure 30: Boxplots of the weighted F1 scores by Random Forest (RF) and Bayesian Neural Network (BNN) models for superficial scab (Scabsuper) disease using environmental predictors.
@@
Figure 31: Boxplots with the weighted F1 scores (y-axis) by Random Forest and Bayesian Neural Network (NN) models (x-axis) for yield per plant. The models including both types of predictors outperform other models, yet models including OTU data alone (OTU-S3) are comparable which suggests that the microbial information indeed contains signal to predict the disease outcome on its own. However, OTU data is more expensive to collect, and perhaps not necessary, given that the model without OTU data (Soil) performs just as accurately (blue dashed line).
@@
Figure 32: Boxplots with the weighted F1 scores (y-axis) by Random Forest and Bayesian Neural Network (NN) models (x-axis) for yield per meter. The models including both types of predictors outperform other models, yet models including OTU data alone (OTU-S3) are comparable which suggests that the microbial information indeed contains signal to predict the disease outcome on its own. However, OTU data is more expensive to collect, and perhaps not necessary, given that the model without OTU data (Soil) performs just as accurately (blue dashed line).
@@
Figure 33: Boxplots with the weighted F1 scores (y-axis) by Random Forest and Bayesian Neural Network (NN) models (x-axis) for black scurf disease. The models including both types of predictors outperform other models, yet models including OTU data alone (OTU-S3) are comparable which suggests that the microbial information indeed contains signal to predict the disease outcome on its own. However, OTU data is more expensive to collect, and perhaps not necessary, given that the model without OTU data (Soil) performs just as accurately (blue dashed line).
@@
Figure 34: Boxplots with the weighted F1 scores (y-axis) by Random Forest and Bayesian Neural Network (NN) models (x-axis) for scab disease. The models including both types of predictors outperform other models, yet models including OTU data alone (OTU-S3) are comparable which suggests that the microbial information indeed contains signal to predict the disease outcome on its own. However, OTU data is more expensive to collect, and perhaps not necessary, given that the model without OTU data (Soil) performs just as accurately (blue dashed line).
@@
Figure 35: Boxplots with the weighted F1 scores (y-axis) by Random Forest and Bayesian Neural Network (NN) models (x-axis) for superficial scab (Scabsuper) disease. The models including both types of predictors outperform other models, yet models including OTU data alone (OTU-S3) are comparable which suggests that the microbial information indeed contains signal to predict the disease outcome on its own. However, OTU data is more expensive to collect, and perhaps not necessary, given that the model without OTU data (Soil) performs just as accurately (blue dashed line).

### v2

@@ Methods
We use the SPRING method (semi-parametric rank-based approach for inference in graphical model) [33] in the NetComi package to construct two microbial networks: one corresponding to label 00 and one corresponding to label $1$. To build these networks, the estimated partial correlations are transformed into dissimilarities via the signed distance metric and the corresponding similarities are then used as edge weights. We then compare these networks in order to identify OTUs that have the greatest difference based on degree values. Figure S5 shows an example of two microbial networks with different graphical structures, and Table S5 in the Supplementary Material lists OTUs ranked by the difference in degree in the two microbial networks for diseased and non-diseased classes of pitted scab response. The OTUs are sorted based on the values of degree difference column and similar to ML strategy top 30% of them are selected as important features.
Combination of ML and Network Strategies in Feature Selection
We define a scoring value for each OTU based on whether they are identified as important by ML strategy ($score=1$), by the network comparison ($score=2$), or both ($score=3$). If the OTU is not identified as important by any strategy, it is denoted $score=0$. Figure S6 displays the scoring values of all OTUs at the Phylum level for all responses.
Model descriptions: random forest and Bayesian neural network
We apply two types of classification models to predict potato disease and yield: RF and Bayesian NN. We consider separately three types of predictors: OTU abundances (5 taxonomic levels and 20 normalization/zero replacement strategies described in Section ‘Data Filtering, Normalization, and Zero Replacement’), environmental predictors such as soil characteristics and microbial population density, and a combination of both types. Table 2 lists all the models we consider in this study.
@@
Table S3 shows the number of predictors included in each model as well as the data choices related to taxonomic level or normalization and zero replacement strategies. For example, using all OTUs (first row), we have five taxonomic levels and 20 normalization+zero replacement strategies each, so in total, we have 100 (20 normalization strategies times 5 taxonomic levels) different OTU datasets. For each of these 100 datasets, the number of predictors (i.e., OTUs) would depend on the taxonomic level being analyzed. For example, at the Phylum level, there might be 42 predictors, while at the Genus level, there could be up to 485 predictors.
Random Forest Model
@@
Here, we set different values for parameters (see Table S6 in the Supplementary Material) and tune them using GridsearchCV to find the optimal values for the RF classifier. GridSearchCV uses a “score” method for evaluating the performance of the cross-validated model on the test set. We use the F1 score to evaluate the performance of the model. The F1 score is the harmonic mean of precision and recall, where the proportional contributions of precision and recall to the F1 score are equal. A F1 score of 1 is the best, while a score of 0 is the worst. Finally, when the parameters for the RF model are tuned, we use 20% of samples to report the performance of the final model. We use the weighted average F1 score to evaluate the performance of the models, which is computed by averaging all the per-class F1 scores while accounting for the number of samples in each class.
Bayesian Neural Network Model
Given that neural network models are intrinsically data-hungry and microbiome applications have relatively small sample sizes, Bayesian NN models provide a suitable alternative for small sample sizes as they provide natural protection against overfitting. This is due to the fact that distributions are considered for the parameters in the model which allow us to marginalize them so that the prediction is based on data points alone [9].
We provide details on the Bayesian NN model formulation next. Let $\mathcal{D}=\{x_{i},y_{i}\}_{i=1}^{n}$ be the data with $n$ samples where $x_{i}$ is a $N$-dimensional feature vector in $\mathbb{R}^{N}$ and $y_{i}\in\mathbb{R}$ is the target response. A Bayesian NN model has $L$ layers with $K_{l}$ neurons in the $l^{th}$ layer. The weight parameters are defined in a set of matrices $\mathcal{W}=\{W_{l}\}_{l=1}^{L}$ where each $W_{l}$ corresponds to the weights for the $l^{th}$ layer, and is of size $(K_{l-1}+1)\times K_{l}$. For an input $x_{i}\in\mathbb{R}^{N}$, the neural network maps it to a response $f(\mathcal{W},\textbf{x})$ by multiplying the input of each layer by the weights and then transforming via an activation function such as rectified linear unit (ReLU): $h(a)=\max(0,a)$.
@@
We use the original implementation of Bayesian NNs denoted “software for flexible Bayesian modeling and Markov chain sampling" [20] for training and testing on our data. For each Bayesian NN model, we use five hidden layers, one input layer and one output layer. The input layer has $n$ input neurons with $n$ being the number of predictors for the model. This number varies depending on the predictors to include. For example, for the ALL-OTU model for the Phylum level, we have 42 predictors (Table 2), and thus the input layer has $N=42$ neurons. For each hidden layer, we have $3N$ hidden neurons and one bias node. For the output layer, we only have one neuron with a binary value for the response. We use the hyperbolic tangent as the activation function: $f(x)=tanh(x)$ as opposed to the sigmoid function, also built in the software, as we found in practice that tanh performs much better.
We choose a hierarchical zero-mean Gaussian prior on all weights and on bias nodes. The prior for the weight from node $i$ of one layer to node $j$ of another layer is formulated as follows [20]: $\mathcal{P}(w_{ij}|\sigma_{w_{i}},\sigma_{a_{j}})=\frac{1}{\sqrt{2\pi}\sigma_{w_{i}}\sigma_{a_{j}}}\exp(-\frac{w_{ij}^{2}}{2\sigma_{w_{i}}^{2}\sigma_{a_{j}}^{2}})$, where $\sigma_{w_{i}}$ are the hyperparameters for the standard deviation of the priors for node $i$ of layer $w$, and $\sigma_{a_{j}}$ is the hyper parameter for the standard deviation of the priors for node $j$ of layer $a$. Let $\tau_{w_{i}}=\sigma_{w_{i}}^{-2}$, then the hyper prior is given by
@@
Given that HMC does not scale well for high dimensional parameter spaces and large datasets [31], we did not fit the Bayesian NN on all the models described in Table 2. In particular, we do not fit a Bayesian NN model on the Genus level for the ALL-OTUs as this model would involve over 9 million weights in the network with 485 input neurons.
Full Model Selection
FMS [25, 26, 24] involves the process of listing all data preprocessing steps, model options and selection of predictors, and using a decision tree model to identify the choices that yield the highest measure of performance. Here, we fit a FMS strategy with the following options: 1) type of normalization, 2) type of zero replacement, 3) taxonomic level, and 4) data augmentation. We combine the type of normalization and type of zero replacement strategy into one variable (denoted $\text{NM}_{i}$ for $i=1,\dots,20$). We focus on the weighted F1 score as measure of performance, and we include all OTU predictors (that is, we do not consider feature selection as one of the options to compare). We build the regression decision tree by using the DecisionTreeRegressor which is available in scikit-learn [21]. We use the default parameters in the DecisionTreeRegressor such as "squared error" as the criterion to measure the quality of a split, a minimum number of 2 for the samples required to split an internal node, and a minimum number of 1 sample required to be at a leaf node. In order to create an informative decision tree that can be interpreted, we use a maximum depth of 4.
@@ Acknowledgements
Level # OTUs # OTUs after filtering
Phylum 57 42
Class 152 108
Order 378 224
Family 476 255
Genus 1319 485
ALL-OTU 5 20 42–485
OTU-S0 5 1 7–54
OTU-S1 5 1 5–134
OTU-S2 5 1 5–134
OTU-S3 5 1 7–54
Alpha 5 6 9
Soil 6 12
DS 6 4
Input: All samples per varietySubset of samples only for Russet varietyCompute the median of the yield valuesFor each sample in Russet varietyYield value $>$ medianLabel = 1Label = 0yesno
@@
Input: training set of original OTUs & all labels# Label 00 $<$ 400Select Label 00# Label $1$ $<$ 400Select Label $1$For each potato varietySelect the subset of samples for that varietyCalculate the mean and variance of each OTU in the subsetGenerate noises $\mathcal{N}\sim(\mu/100,\sigma/100)$Augmented value = max(0, original+noise)Add this augmented sampleOutput: 800 samplesyesnoyesno
@@
Figure 3: Boxplots of weighted F1 scores for random forest models for different types of predictors (rows) and different yield or disease outcomes (columns). For description of the rows, see Table 2 in the main text. The range of each boxplot depicts the weighted F1 scores for datasets in different taxonomic levels and different normalization and zero replacement strategies. The dashed line corresponds to the weighted F1 score when fitting the model with random datasets(see Section A).
Figure 4: Boxplots of weighted F1 scores for Bayesian neural network models for different types of predictors (rows) and different yield or disease outcomes (columns). For description of the rows, see Table 2 in the main text. The range of each boxplot depicts the weighted F1 scores for datasets in different taxonomic levels and different normalization and zero replacement strategies. The dashed line corresponds to the weighted F1 score when fitting the model with random datasets (see Section A).
@@
Appendix A More power to predict pitted scab disease in microbiome data compared to random data
For the pitted scab on the top row, the F1 scores for the original dataset (red dots) are on the upper right quadrant compared to points from random datasets (blue dots). This means that the F1 scores on the original data are higher than F1 scores on random data, and this, the original data has more predictive power than random data. However, for the superficial scab on the bottom row, many random datasets show better performance (higher F1 score values) than the real data, and thus, the real data does not contain much information to predict this disease. Other diseases as well as yield outcomes also underperform compared to random data as shown in Figure 8 in the Supplementary Material. We highlight that pitted scab is the most severe case of the potato disease common scab, and possibly a stronger signal of the soil microbial community composition and corresponding disease-suppressiveness than other measures of the disease such as superficial scab.
A.1 Predictive Power of Microbiome Data: Comparison to Random Data
Strategy 1:
Strategy 2:
Strategy 3:
Real microbiome matrix is used with a permuted response vector.
Strategy 4:
For each strategy, we generate $N=200$ random datasets to train the RF model. Let $F_{i}$ for $i=1,...,N$, denote the weighted F1 scores for the randomized data and let $F_{original}$ be the weighted F1 score on the real data. Let $X=\sum_{i=1}^{N}\mathbb{I}(F_{i}>F_{original})$ be the number of random datasets that perform better than the read data. We use the exceeding value (EV) $EV=\frac{X}{N}$ as test statistic to test the null hypothesis that the real data performs just as random data in the prediction on disease outcomes. The EV denotes the percentage of random datasets that perform better than the real microbiome data out of $N$ random datasets. If $EV<\alpha$, we reject the null hypothesis and conclude that there is predictive power (beyond random) on the real microbiome data. Here, we set $\alpha=0.05$.
@@ A.1 Predictive Power of Microbiome Data: Comparison to Random Data
Figure 30: Boxplots of the weighted F1 scores by random forest (RF) and Bayesian neural network (Bayesian NN) models for pitted scab disease using environmental predictors (Soil: Soil chemistry, DS: Microbial population density in soil).
@@
Figure 31: Boxplots of the weighted F1 scores by Random Forest (RF) and Bayesian Neural Network (BNN) models for yield by plant using environmental predictors (Soil: Soil chemistry, DS: Microbial population density in soil).
Figure 32: Boxplots of the weighted F1 scores by Random Forest (RF) and Bayesian Neural Network (BNN) models for yield by meter using environmental predictors (Soil: Soil chemistry, DS: Microbial population density in soil).
Figure 33: Boxplots of the weighted F1 scores by Random Forest (RF) and Bayesian Neural Network (BNN) models for black scurf disease using environmental predictors (Soil: Soil chemistry, DS: Microbial population density in soil).
Figure 34: Boxplots of the weighted F1 scores by Random Forest (RF) and Bayesian Neural Network (BNN) models for scab disease outcome using environmental predictors (Soil: Soil chemistry, DS: Microbial population density in soil).
Figure 35: Boxplots of the weighted F1 scores by Random Forest (RF) and Bayesian Neural Network (BNN) models for superficial scab (Scabsuper) disease using environmental predictors (Soil: Soil chemistry, DS: Microbial population density in soil).
@@
Figure 36: Boxplots with the weighted F1 scores (y-axis) by Random Forest and Bayesian Neural Network (NN) models for yield per plant. The models including both types of predictors outperform other models, yet models including OTU data alone (OTU-S3) are comparable which suggests that the microbial information indeed contains signal to predict the disease outcome on its own. However, OTU data is more expensive to collect, and perhaps not necessary, given that the model without OTU data (Soil) performs just as accurately (blue dashed line).
@@
Figure 37: Boxplots with the weighted F1 scores (y-axis) by Random Forest and Bayesian Neural Network (NN) models for yield per meter. The models including both types of predictors outperform other models, yet models including OTU data alone (OTU-S3) are comparable which suggests that the microbial information indeed contains signal to predict the disease outcome on its own. However, OTU data is more expensive to collect, and perhaps not necessary, given that the model without OTU data (Soil) performs just as accurately (blue dashed line).
@@
Figure 38: Boxplots with the weighted F1 scores (y-axis) by Random Forest and Bayesian Neural Network (NN) models for black scurf disease. The models including both types of predictors outperform other models, yet models including OTU data alone (OTU-S3) are comparable which suggests that the microbial information indeed contains signal to predict the disease outcome on its own. However, OTU data is more expensive to collect, and perhaps not necessary, given that the model without OTU data (Soil) performs just as accurately (blue dashed line).
@@
Figure 39: Boxplots with the weighted F1 scores (y-axis) by Random Forest and Bayesian Neural Network (NN) models for scab disease. The models including both types of predictors outperform other models, yet models including OTU data alone (OTU-S3) are comparable which suggests that the microbial information indeed contains signal to predict the disease outcome on its own. However, OTU data is more expensive to collect, and perhaps not necessary, given that the model without OTU data (Soil) performs just as accurately (blue dashed line).
@@
Figure 40: Boxplots with the weighted F1 scores (y-axis) by Random Forest and Bayesian Neural Network (NN) models for superficial scab (Scabsuper) disease. The models including both types of predictors outperform other models, yet models including OTU data alone (OTU-S3) are comparable which suggests that the microbial information indeed contains signal to predict the disease outcome on its own. However, OTU data is more expensive to collect, and perhaps not necessary, given that the model without OTU data (Soil) performs just as accurately (blue dashed line).